# KeyDiff: Key Similarity-Based KV Cache Eviction for Long-Context LLM Inference in Resource-Constrained Environments

### v2

@@ Abstract
In this work, we demonstrate that distinctive keys during LLM inference tend to have high attention scores. We explore this phenomenon and propose KeyDiff, a training-free KV cache eviction method based on key similarity. This method facilitates the deployment of LLM-based application requiring long input prompts in resource-constrained environments with limited memory and compute budgets. Unlike other KV cache eviction methods, KeyDiff can process arbitrarily long prompts within strict resource constraints and efficiently generate responses. We demonstrate that KeyDiff computes the optimal solution to a KV cache selection problem that maximizes key diversity, providing a theoretical understanding of KeyDiff. Notably, KeyDiff does not rely on attention scores, allowing the use of optimized attention mechanisms like FlashAttention. We demonstrate the effectiveness of KeyDiff across diverse tasks and models, illustrating a performance gap of less than 0.04% with 8K cache budget ($\sim$ 23% KV cache reduction) from the non-evicting baseline on the LongBench benchmark for Llama 3.1-8B and Llama 3.2-3B.
Machine Learning, ICML
@@ 1 Introduction
Large language models (LLMs) have revolutionized natural language processing, enabling tasks like document summarization, code generation, question answering (Brown et al., 2020; Raffel et al., 2020; Touvron et al., 2023; Dubey et al., 2024). Many exciting applications of LLMs, such as retrieval augmented generation (Lewis et al., 2020) and reasoning tasks (Wei et al., 2022; Kojima et al., 2022; Yao et al., 2024), require very long input prompts, or long context support, to provide a valuable end-user experience. However, deploying these models in production environments has grown in importance and managing long input prompts remains a challenge due to large compute and memory requirements.
Due to cost and privacy concerns, there has been a recent focus on deploying LLMs on the edge, i.e. directly on devices such as mobile phones or personal computers (Alizadeh et al., 2023; Liu et al., 2024c; van Baalen et al., 2024; Xu et al., 2024a). This imposes tight resource constraints, such as limited or shared compute resources, limited available DRAM, and even power limits in the mobile setting. Many techniques, such as system optimization (Alizadeh et al., 2023; Liu et al., 2024c; Xu et al., 2024a), model quantization (van Baalen et al., 2024), and efficient text generation (Goel et al., 2024) help to enable on-device LLM inference, but the main challenge for long context application on-device deployment is effectively managing key and value states.
In a naïve implementation of autoregressive text generation, key and value states (KVs) must be recomputed for each new generated token. In a decoder-only transformer model, one can instead store the previously computed KVs in a KV cache for later use, allowing for a significant decrease in LLM inference latency. This comes at the cost of maintaining the KV cache, whose memory footprint grows linearly with the input sequence length (Yang et al., 2024a) and will violate any imposed memory limit with a sufficiently large input.
One solution to bound this memory overhead is to fix the number of KVs in the cache, or cache budget, and use a cache eviction policy to remove KVs from the cache when the budget has been exceeded, bounding the KV cache memory usage. However, state-of-the-art cache eviction policies retain KVs with high attention weights (Xiao et al., 2024; Oren et al., 2024; Zhang et al., 2024b), which requires explicitly materializing the attention matrix of the entire input prompt. This results in large intermediate memory overhead during cache eviction, often violating memory limits.
This intermediate cache memory overhead can be mitigated by segmenting the input prompt into blocks, sequentially processing each block by the LLM and caching the intermediate KVs (Kwon et al., 2023; Agrawal et al., 2023; Xu et al., 2024b). The cache eviction policy can be applied block-wise to satisfy imposed memory constraints (Figure 1). However, our experiments show that applying existing attention-based eviction methods with block prompt processing results in performance degradation of downstream tasks (Table 1). This is due to the limited scope of the attention computation during block prompt processing: attention scores are only computed for a subsequence of the original prompt and only contain local information. These partial scores may not be representative of a token’s global importance across the entire prompt and can result in incorrect cache eviction. This means that we need an eviction method that is independent from attention and cheap to compute during inference while maintaining model capability.
The “attention sink” phenomenon (Xiao et al., 2024; Sun et al., 2024) describes the observation that, in decoder-only transformer-based LLMs, high attention scores are assigned to the first few tokens in the KV cache regardless of the input. This implies that eviction can be determined by only inspecting KVs and that an attention-free cache eviction criteria exists. In fact, we observe that keys with lower average pairwise cosine similarity also have higher attention outputs, as shown in Figure 2.
Motivated by these observations, we propose an eviction method called KeyDiff, that evicts keys from the cache with the highest cosine similarity to a cache representative called an anchor. We also show that KeyDiff computes the optimal solution of the KV cache selection problem that minimizes pairwise key cosine similarity. Our experiments demonstrate that KeyDiff can achieve an accuracy drop of $\leq$ .04% on LongBench (Bai et al., 2024) with 8K token budget ($\sim$ 23% KV cache reduction), and can match or outperform state-of-the-art eviction methods, especially when using block prompt processing. Our contributions are as follows:
  • We present the empirical observation that low pairwise cosine similarity among keys is correlated with higher attention output.
  • We design an eviction policy based on this observation, called KeyDiff, that enables long context LLM inference with limited memory overhead.
  • We show that KeyDiff computes the optimal solution to a KV cache selection problem that maximizes key diversity measured by pairwise cosine similarity.
  • We demonstrate the effectiveness of KeyDiff on the LongBench task suite, achieving only 1.5% and 0.04% performance gap from the eviction-free baseline, using 6K and 8K cache budgets, respectively.
@@ 2.1 Transformers
The Transformer architecture (Vaswani, 2017) processes input data using a sequence of transformer blocks. A transformer block $f$ takes a sequence $X=(x_{1},x_{2},\ldots,x_{T})\in\mathbb{R}^{T\times d}$ as input and applies the causal self-attention operator $\operatorname{Attention}$ followed by a feed-forward network $\operatorname{FF}$ with optional gating (Shazeer, 2020) to produce the output $X^{\prime}=(x^{\prime}_{1},x^{\prime}_{2},\ldots,x^{\prime}_{T})\in\mathbb{R}^{T\times d}$:
@@
$O^{\operatorname{attn}}=\operatorname{Softmax}\left(\frac{QK^{\top}}{\sqrt{d}}+M\right)V=AV$ (2)
@@ 2.2 KV Caching
$o^{\operatorname{attn}}_{T+1}=\operatorname{Softmax}\left(\frac{q_{T+1}[K\|k_{T+1}]^{\top}}{\sqrt{d}}+M\right)[V\|v_{T+1}],$ (3)
@@
KV caching dramatically reduces the latency of $\operatorname{Attention}$ by only computing $k_{T+1}$, $q_{T+1}$, $v_{T+1}$ for each token $x_{T+1}$ and reusing the KVs in $\mathcal{C}$. However, the size of the KV cache increases linearly with the number of processed tokens and dominates the memory footprint in long-context applications (Yang et al., 2024a).
@@ 2.3 KV Cache Eviction Methods
$K^{\prime}$ $=\texttt{gather}(K,S)$
$V^{\prime}$ $=\texttt{gather}(V,S)$
@@ 2.4 KV Caching in Resource-Constrained Environments
Existing eviction policies like Zhang et al. (2024b); Oren et al. (2024) focus on processing the entire input prompt at once: KVs are computed for each token in the prompt and stored in a cache $\mathcal{C}$, then the eviction policy $\pi_{N}$ is applied to reduce the number of tokens in $\mathcal{C}^{\prime}$ to $N$, before token generation. However, the intermediate cache $\mathcal{C}$ before eviction will grow to the size of the input prompt. This can often exceed model’s allocated memory limit when deploying long context applications in resource constrained environments.
As demonstrated in efficient LLM inference frameworks (Agrawal et al., 2023; Holmes et al., 2024; Kwon et al., 2023; Xu et al., 2024a), one solution is to apply $\pi_{N}$ more frequently by segmenting $X$ into non-overlapping blocks $X=[X_{0},X_{1},\ldots,X_{m-1}]$, where $X_{i}=[x_{Bi},\ldots,x_{B(i+1)-1}]$, $B$ is the block size, and $m=\lceil T/B\rceil$, and iteratively updating the cache by exploiting causality, applying Equation 4 in a block-wise fashion:
@@
where $k_{Bi:B(i+1)-1}$ and $v_{Bi:B(i+1)-1}$ are the keys and values selected from $X_{i}$ respectively, and $\mathcal{C}_{i}=(K_{i},V_{i})$ is the KV cache after processing the first $i$ prompt blocks. As in Equation 4, we concatenate the $B$ new KVs to the current cache, apply the eviction policy $\pi_{N}$ and update the cache in Equation 6.
@@ 3 Method
We demonstrate a negative correlation between attention scores and the similarity among keys (Section 3.1) and leverage this observation to develop KeyDiff (Section 3.2), followed by a theoretical justification of the method and preliminary evidence of its efficacy (Section 3.3).
@@ 3.1 Correlation of Attention Scores and Key Dissimilarity
To address the shortcomings of attention-based KV cache eviction in Section 2.4, we develop an alternative attention-free scoring metric that retains significant KVs across blocks while being resource efficient. We recall the “attention sink” phenomenon: LLMs often assign high attention weight to the first few tokens, regardless of the input (Xiao et al., 2024; Sun et al., 2024); these highly weighted tokens are called sink tokens. However, the index of the sink tokens can vary across heads and layers and reside deeper in the sequence than the first few tokens. This observation motivates the following hypothesis: high attention scores can be determined by the intrinsic properties of the keys rather than by any particular combination of keys and queries.
@@
We evaluate our hypothesis by inspecting the cosine similarities between keys computed inside an attention block. We visualize the pairwise cosine similarities between keys along with the attention weights in two particular heads and layers in Figure 2. We observe that keys with lower cosine similarity with other keys exhibit higher relative attention scores regardless of the choice of query, such as the 4th and 15th keys in Figure 2(a), or the 1st key in Figure 2(b). Pairwise cosine similarity of keys is solely a function of the keys in the cache, which are independent of input queries; the surprising aspect of Figure 2 is the negative correlation with attention weights. These distinctive keys essentially recover the attention sink phenomenon (Xiao et al., 2024).
@@ 3.2 KeyDiff
$S$ $=\texttt{topk}(-\operatorname{CosSim}(K)\mathbf{1},N)$ (7)
$K^{\prime}$ $=\texttt{gather}(K,S)$
$V^{\prime}$ $=\texttt{gather}(V,S)$
where $K\in\mathbb{R}^{n\times d}$ and $V\in\mathbb{R}^{n\times d}$ are the cached keys and values, $\operatorname{CosSim}(K)\in\mathbb{R}^{n\times n}$ is the pairwise cosine similarity matrix of keys in $K$ with $\operatorname{CosSim}(K)_{ij}=\frac{\langle k_{i},k_{j}\rangle}{\|k_{i}\|\|k_{j}\|}$, and $\mathbf{1}\in\mathbb{R}^{n}$ is a vector of ones.
@@
Unlike attention-based eviction policies, KeyDiff does not require access to the attention weights $A$, facilitating optimized attention kernels that do not materialize $A$, such as FlashAttention (Dao et al., 2022). However, computing the pairwise cosine similarities runs in $\mathcal{O}(n^{2})$ time. Fortunately, the operation
$\texttt{topk}(-\operatorname{CosSim}(K)\mathbf{1},N)$
can be computed in $\mathcal{O}(n)$ operations on average by computing the score of each token as
$s_{i}=-\operatorname{CosSim}(\mu(\hat{K}),\hat{k}_{i})$
where $\mu(\hat{K})=\frac{1}{n}\sum_{i=1}^{n}\hat{k}_{i}$ and $\hat{k}_{i}=\frac{k_{i}}{||k_{i}||}$, and retaining the tokens with the $N$ largest scores with a variation of the quick select algorithm. This simplification is derived in Section 3.3.
The eviction score of the efficient expression of KeyDiff is defined as follows:
$s$ $=-\operatorname{CosSim}(\mu(\hat{K}),K)$ (8)
$S$ $=\texttt{topk}(s,N)$
We refer to $\mu(\hat{K})$ as the anchor vector. Experimentation has shown that the anchor vector $\mu(\hat{K})$ can be replaced with $\mu(K)$ without losing accuracy (see Table 2). We evaluate this efficient version of KeyDiff described in Figure 3 in the following sections.
3.3 Derivation of KeyDiff from an Optimization Perspective
[…]
This is a combinatorial optimization problem, which is difficult to solve efficiently, particularly during inference. However, we can relax Section 3.3 to produce a more tractable approximation to the original problem. First, we rewrite Section 3.3 by normalizing keys $k_{i}$ such that $\hat{k}_{i}=\frac{k_{i}}{\|k_{i}\|}$ resulting in:
[…]
Key Cache Diversity and KeyDiff
To empirically verify KeyDiff’s ability to retain diverse keys, we apply PCA to the keys computed in an attention block of Llama 3.2-3B-Instruct after evaluating a long context prompt. We visualize the distribution of retained and evicted keys by applying sink attention (Xiao et al., 2024), TOVA (Oren et al., 2024), and KeyDiff as eviction policies in Figure 4. Visual observation reveals the tokens retained by sink attention and TOVA tend to tightly cluster together while KeyDiff’s retained tokens are more evenly distributed. This observation generalizes across heads and layers, as shown in Section A.3.
We also visualize the log determinant of the Gram matrix of the key cache $\log(\det(KK^{T}))$ generated using different eviction policies in Figure 5. This quantity corresponds to the volume of space spanned by the keys in $\mathcal{C}$. The distribution of volumes for KeyDiff attains higher values, indicating that the retained keys are more distinctive than TOVA and sink, corroborating the results of Figure 4. Details of how these plots were generated are discussed in Section B.1.

### v3

@@ Abstract
We demonstrate that geometrically distinctive keys during LLM inference tend to have high attention scores. Based on the phenomenon we propose KeyDiff, a training-free KV cache eviction method based solely on key similarity. Unlike other KV cache eviction methods, KeyDiff can process arbitrarily long prompts within strict resource constraints and efficiently generate responses. We provide a theoretical basis for KeyDiff by relating key diversity with attention scores. These results imply KeyDiff can efficiently identify the most important tokens to retain. Notably KeyDiff does not rely on attention scores, allowing the use of optimized attention mechanisms like FlashAttention. Under a strict memory allowance, we demonstrate the effectiveness of KeyDiff for the Llama and Qwen model families by observing a performance gap of less than 0.04% with 8K cache budget ($\sim$ 23% KV cache reduction) from the non-evicting baseline on LongBench for Llama 3.1-8B and Llama 3.2-3B. We also observe near baseline performance for Deepseek-R1-Distill-Llama-8B on the Math500 reasoning benchmark and decrease end-to-end inference latency by up to 30% compared to the other token-eviction methods.
@@ 1 Introduction
Key-Value (KV) caching is a standard technique to accelerate large language model (LLM) inference that reuses key and value states (KVs) from previously processed tokens, enabling efficient autoregressive generation. This is crucial for long-context applications such as document summarization, code generation, question answering [7, 25, 31, 10], retrieval augmented generation [19] and reasoning [34, 17, 41]. However, the memory footprint of the stored KV cache grows linearly with input length, which becomes a bottleneck in memory-constrained environments.
This challenge is particularly acute for LLM inference on edge device, where compute, memory, and power resources are limited [3, 22, 32, 36]. While KV cache eviction policies have been proposed to bound memory overhead by removing unimportant KVs (often measured by attention scores) [35, 24, 43], they typically process the entire prompt at once and violate memory constraints during intermediate computation.
To enforce strict memory bounds throughout the prompt prefill and token generation inference phases, we adopt a block-wise inference strategy: the input prompt is divided into smaller blocks which are processed sequentially by the model, similar to [18, 1, 37]. After processing each block, we evict some of the cached KVs by according to an eviction policy that scores each KV, as illustrated in Figure 1. Unlike previous approaches that apply eviction after processing the entire prompt, this strategy satisfies memory constraints throughout the full inference process. However, we observe a degradation in accuracy when applying existing eviction methods in this setting (Table 1).
We hypothesize that the performance drop stems from a mismatch in design: existing eviction methods assume access to full-prompt attention, where key importance is computed over the entire input. During block prompt processing, however, attention is computed using only the current block’s tokens without access to future blocks. As a result, attention scores based on a limited context often fail to reflect a token’s true importance across the full prompt.
To this end, we observe that keys with lower average pairwise cosine similarity tend to receive higher attention scores across a variety of inputs. This suggests that key diversity serves as a strong proxy for global token importance, even without access to future tokens. This insight enables an attention-free approach to cache eviction that is based on the geometry of the cached keys.
Motivated by these observations, we propose KeyDiff, an attention-free cache eviction method that removes redundancy among cached keys, operates effectively during block-wise inference, and avoids excessive memory overhead. Our contributions are summarized as follows:
  • Insight. We observe that lower pairwise cosine similarity among keys correlates with higher attention scores, suggesting its utility as a proxy for token importance. (Section 3.1)
  • Method. We introduce KeyDiff, an eviction strategy that selects keys based on their similarity to other cached entries without relying on attention scores or future tokens. (Section 3.2)
  • Theory. Through our analysis of key and query geometry we provide a theoretical understanding how/why KeyDiff works. We also show that KeyDiff solves an optimal subset selection problem that maximizes key diversity. (Section 3.3)
  • Performance. KeyDiff achieves $\leq 1.5\%$ and $\leq 0.04\%$ accuracy drop on LongBench with 6K and 8K cache budgets, respectively, outperforming state-of-the-art eviction methods across Llama and Qwen models (Section 4.2), and near non-evicting baseline performance for Deepseek-R1-Distill-Llama-8B on the Math-500 reasoning benchmark. (Section 4.3)
  • Efficiency. We observe up to $30\%$ end-to-end inference latency reduction using KeyDiff compared to existing KV cache eviction methods. (Section 4.5)
@@ 2.1 Transformers
The Transformer architecture [33] processes input data using a sequence of transformer blocks. A transformer block $f$ takes a sequence $X=(x_{1},x_{2},\ldots,x_{T})\in\mathbb{R}^{T\times d}$ as input and applies the causal self-attention operator $\operatorname{Attention}$ followed by a feed-forward network $\operatorname{FF}$ with optional gating [28] to produce the output $X^{\prime}=(x^{\prime}_{1},x^{\prime}_{2},\ldots,x^{\prime}_{T})\in\mathbb{R}^{T\times d}$:
@@
$O^{\operatorname{attn}}=\operatorname{Softmax}\left(\Large\nicefrac{{QK^{\top}}}{{\sqrt{d}}}+M\right)V=AV$ (2)
@@ 2.2 KV Caching
$o^{\operatorname{attn}}_{T+1}=\operatorname{Softmax}\left(\Large\nicefrac{{q_{T+1}[K\|k_{T+1}]^{\top}}}{{\sqrt{d}}}+M\right)[V\|v_{T+1}],$ (3)
@@
KV caching dramatically reduces the latency of $\operatorname{Attention}$ by only computing $k_{T+1}$, $q_{T+1}$, $v_{T+1}$ for each token $x_{T+1}$ and reusing the KVs in $\mathcal{C}$. However, the size of the KV cache increases linearly with the number of processed tokens and dominates the memory footprint in long-context applications [39].
@@ 2.3 KV Cache Eviction Methods
$K^{\prime}$ $=\texttt{gather}(K,S),\quad V^{\prime}=\texttt{gather}(V,S)$
@@ 2.4 KV Caching in Resource-Constrained Environments
Existing eviction policies like Zhang et al. [43], Oren et al. [24] focus on processing the entire input prompt at once: KVs are computed for each token in the prompt and stored in a cache $\mathcal{C}$, then the eviction policy $\pi_{N}$ is applied to reduce the number of tokens in $\mathcal{C}^{\prime}$ to $N$, before token generation. However, the intermediate cache $\mathcal{C}$ before eviction will grow to the size of the input prompt. This can often exceed model’s allocated memory limit when deploying long context applications in resource constrained environments.
As demonstrated in efficient LLM inference frameworks [1, 14, 18, 36], one solution is to apply $\pi_{N}$ more frequently by segmenting $X$ into non-overlapping blocks $X=[X_{0},X_{1},\ldots,X_{m-1}]$, where $X_{i}=[x_{Bi},\ldots,x_{B(i+1)-1}]$, $B$ is the block size, and $m=\lceil T/B\rceil$, and iteratively updating the cache by exploiting causality, applying Equation 4 in a block-wise fashion:
@@
where $k_{Bi:B(i+1)-1}$ and $v_{Bi:B(i+1)-1}$ are the keys and values selected from $X_{i}$ respectively, and $\mathcal{C}_{i}=(K_{i},V_{i})$ is the KV cache after processing the first $i$ prompt blocks. As in Equation 4, we concatenate the $B$ new KVs to the current cache, apply $\pi_{N}$ and update the cache in Equation 6.
@@ 3 Method
We demonstrate a negative correlation between attention scores and the cosine similarity among keys (Section 3.1) and leverage this observation to develop KeyDiff (Section 3.2), followed by a theoretical justification of KeyDiff (Section 3.3) and preliminary evidence of its efficacy (Section C.2).
@@ 3.1 Correlation of Attention Scores and Key Dissimilarity
To address the shortcomings of attention-based KV cache eviction in Section 2.4, we develop an alternative attention-free scoring metric that retains significant KVs across blocks while being resource efficient. We recall the “attention sink" phenomenon: LLMs often assign high attention weight to the first few tokens, regardless of the input [35, 29]; these highly weighted tokens are called sink tokens. However, the index of the sink tokens can vary across heads and layers and reside deeper in the sequence than the first few tokens. This observation motivates the following hypothesis: high attention scores can be determined by the intrinsic properties of the keys rather than by any particular combination of keys and queries.
@@
We evaluate our hypothesis by inspecting the cosine similarities between keys computed inside an attention block. We visualize the pairwise cosine similarities between keys along with the attention weights in two particular heads and layers in Figure 2. We observe that keys with lower cosine similarity with other keys exhibit higher relative attention scores regardless of the choice of query, such as the 4th and 15th keys in Figure 2(a), or the 1st key in Figure 2(b). Pairwise cosine similarity of keys is solely a function of the keys in the cache, which are independent of input queries; the surprising aspect of Figure 2 is the negative correlation with attention weights. These distinctive keys essentially recover the attention sink phenomenon [35].
@@ 3.2 KeyDiff
$S$ $=\texttt{topk}(-\operatorname{CosSim}(K)\mathbf{1},N),$ (7)
$K^{\prime}$ $=\texttt{gather}(K,S),\quad V^{\prime}=\texttt{gather}(V,S)$
where $K\in\mathbb{R}^{n\times d}$ and $V\in\mathbb{R}^{n\times d}$ are the cached keys and values, $\operatorname{CosSim}(K)\in\mathbb{R}^{n\times n}$ is the pairwise cosine similarity matrix of keys in $K$ with $\operatorname{CosSim}(K)_{ij}=\frac{k_{i}\cdot k_{j}}{\|k_{i}\|\|k_{j}\|}$, and $\mathbf{1}\in\mathbb{R}^{n}$ is a vector of ones.
@@
Unlike attention-based eviction policies, KeyDiff does not require access to the attention weights $A$, facilitating optimized attention kernels that do not materialize $A$ such as FlashAttention [8]. However, computing the pairwise cosine similarities runs in $\mathcal{O}(n^{2})$ time. Fortunately, we can compute the score of each token in Equation 7 in $\mathcal{O}(n)$ as follows:
$S$ $=\texttt{topk}(-\operatorname{CosSim}(\mu(\hat{K}),\hat{k}_{i}),N)$ (8)
where $\mu(\hat{K})=\frac{1}{n}\sum_{i=1}^{n}\hat{k}_{i}$ and $\hat{k}_{i}=\frac{k_{i}}{||k_{i}||}$. We refer to $\mu(\hat{K})$ as the anchor vector. We show this formulation retains the same KVs of Equation 7 under a mild condition. (see Section C.2). Our experimentation has shown that the anchor vector $\mu(\hat{K})$ can be replaced with $\mu(K)$ without losing accuracy (see Table 14). We evaluate the efficient KeyDiff described in Figure 3 using unnormalized keys $k$ in all subsequent sections. Figures 5, 8, 9 and 10 visualize the keys retained and evicted by sink attention [35], TOVA [24] and KeyDiff via PCA. KeyDiff retains more varied keys. A full complexity and FLOP analysis can be found in Appendices B and B.1
KeyDiff with Sliding Window
In tasks such as reasoning and coding, where the most recent tokens are often important, we can augment KeyDiff and its efficient variant to use a percentage of the cache budget for a sliding window [6], which we call KeyDiff with sliding window. This extension introduces no complexity or memory overhead and we observe better results on certain tasks than vanilla KeyDiff (Tables 13 and E).
3.3 Why KeyDiff Works: A Theoretical Perspective
To solidify the theoretical foundation of KeyDiff and show that KeyDiff ultimately selects keys most aligned with queries, we prove the following two results. We first validate the relationship between cosine similarity and attention scores observed in Figure 2 by bounding the attention score of a new incoming key $k^{\star}$ in terms cosine similarity with a fixed query $q$:
Lemma 3.1.
Suppose that for a fixed query token $q$, there is a set of key tokens $\{k_{i}\}_{i=1}^{n}$ such that $||k_{i}||_{2}^{2}<M,\;\forall\;i$. Without loss of generality suppose $||q||=1$ and assume ${k^{*}}$ is a key not in $\{k_{i}\}_{i=1}^{n}$ with $||{k^{*}}||_{2}^{2}<M$ that has attention weight $w>0$. Then, for $n\to\infty$,
$\frac{-\log(1-w)}{2M}-1\leq\textrm{CosSim}({k^{*}},q)$
We then establish a relationship between the cosine similarities of $k^{\star}$, $q$, and the mean of prior keys $\bar{k}$:
Theorem 3.2.
Consider tokens ${k^{*}}$, $q$ as above, and the average of the keys tokens $\bar{k}$. Suppose $\textrm{CosSim}({k^{*}},q)=\beta_{q}>0$ and $\textrm{CosSim}(\bar{k},q)=\alpha_{q}<0$. Then
$\textrm{CosSim}(\bar{k},{k^{*}})\leq 1+\alpha_{q}\beta_{q}-0.5\alpha_{q}^{2}-0.5\beta_{q}^{2}.$ (9)
Figure 4: PCA embedding of keys and queries from Llama 3.2 3B
By combining Lemmas 3.1 and 3.2, we establish a relationship between the attention weight $w$ and the KeyDiff score $\mathrm{CosSim}(\bar{k},k^{*})$. As $\mathrm{CosSim}(\bar{k},q)$ decreases and $\mathrm{CosSim}(k^{*},q)$ increases (along with the attention weight $w$), then $\mathrm{CosSim}(\bar{k},k^{*})$ tends to $-1$: this means KeyDiff selects distinct keys most aligned with $q$. We visualize this in Figure 4 with a PCA embedding of keys and queries from a single head of Llama 3.2 3B, highlighting the relationship between top scoring keys via KeyDiff, the anchor vector and queries. Similar trends are found from the other layers and heads as shown in Figure 11. The proofs of Lemmas 3.1 and 3.2 are in Section C.3, along with empirical motivation for the chosen assumptions.